# Can Slow-Thinking LLMs Reason Over Time? Empirical Studies in Time Series Forecasting

### v3

@@ Abstract.
Time series forecasting (TSF) is a fundamental task, with methods ranging from classical statistical models to modern deep learning and multimodal language models. Despite their effectiveness, most approaches follow a fast-thinking paradigm—focusing on pattern extraction and direct value mapping—while overlooking explicit reasoning over temporal dynamics and contextual semantics, which is crucial for capturing long-range dependencies and regime shifts in real-world data. In contrast, emerging slow-thinking large language models (LLMs), such as OpenAI o1 and DeepSeek-R1, show strong multi-step reasoning abilities across domains, offering an opportunity to reframe TSF as a reasoning-driven generation task. This raises a key question: can slow-thinking LLMs reason effectively over dynamic temporal patterns to support accurate TSF, even in a training-free setting? To answer this, we present TimeReasoner, an empirical study that formulates TSF as a conditional reasoning task. We design prompting strategies, including chain-of-thought and context-augmented reasoning, to elicit inference-time reasoning from pretrained slow-thinking LLMs, and evaluate them on diverse TSF benchmarks across domains and horizons. Our results show that slow-thinking LLMs achieve non-trivial training-free forecasting performance and can outperform prior baselines in most cases. While preliminary, this work offers insights into the reasoning behaviors of slow-thinking LLMs in temporal domains, highlighting both potential and limitations, and pointing toward more interpretable, generalizable, and cross-domain TSF frameworks [^1] https://github.com/realwangjiahao/TimeReasoner.
@@ 1. Introduction
Time series forecasting (TSF) is a fundamental data mining task with wide-ranging applications in domains such as finance (tay2001application), energy, healthcare, and transportation (deb2017review), where accurate forecasts are critical for strategic planning and effective decision-making under uncertainty (box1968some; pan2025onecast). The goal of TSF is to predict future values of target variables by leveraging historical observations together with contextual information, including temporal covariates, static attributes, and exogenous factors (hu2016building; wang2024tabletime; jintaozhang).
@@
Over the past decades, numerous efforts have been proposed (cheng2025comprehensive; NEURIPS2023_2e19dab9), broadly categorized into traditional statistical modeling (lim2021time) and modern data-driven approaches. Traditional statistical models, such as ARIMA (zhang2003time) and Holt-Winters (chatfield1978holt), rely on strong assumptions of linearity, stationarity, and seasonality. While offering clear interpretability through explicit parametric structures, they often struggle with complex, nonlinear, or nonstationary temporal patterns. In contrast, data-driven approaches, particularly deep learning-based models, capture intricate patterns such as trends and seasonality via nonlinear architectures, achieving strong performance across benchmarks (benidis2022deep; zhang2024trajectory) and further improvements in various TSF scenarios (NEURIPS2024_053ee34c; ilbert2024analysing). Recently, with the rapid progress of LLMs, TSF has been explored from multimodal understanding and generative modeling perspectives, leveraging LLMs’ capacity to incorporate heterogeneous contextual features (jiang2024empowering). Follow-up studies extend this direction through large-scale LLM integration and generation-oriented architectures (zhang2024large). However, most existing attempts treat LLMs as black-box sequence mappers, without explicitly exploiting their potential for step-wise temporal reasoning—an ability crucial for long-horizon and context-aware forecasting tasks (tan2024language; gat2023faithful).
@@
Building on this observation, we argue that most TSF methods follow a fast-thinking paradigm (jha2016comparative), mapping historical inputs to future targets through single-step inference with little or no explicit intermediate reasoning (chow2024towards). In this paradigm, models tend to capture short-term temporal dependencies but struggle to integrate heterogeneous contextual information via step-by-step reasoning (kauppinen2007modeling). In contrast, slow-thinking LLMs such as ChatGPT-o1 (o1) and DeepSeek-R1 (guo2025deepseek) have demonstrated strong multi-step reasoning abilities in domains like mathematics, typically by generating intermediate steps before final answers (muennighoff2025s1), a process potentially beneficial for handling long-range dependencies and contextual regime shifts in TSF. However, it still remains under-explored whether such reasoning abilities can be effectively leveraged for time series forecasting, especially in a training-free setting without task-specific fine-tuning.
@@
Clearly, these questions—centered on the potential of slow-thinking LLMs for TSF—are central to rethinking and reformulating TSF. In this work, we explore whether slow-thinking LLMs can serve as forecasters through explicit, step-by-step reasoning. To this end, we reformulate TSF as a training-free conditional reasoning task, where predictions are guided by historical observations and heterogeneous contextual features. Building on the formulation in Figure 1, we propose TimeReasoner, a multimodal inference-time reasoning framework that leverages LLMs’ general reasoning capabilities to produce both interpretable forecasts and structured reasoning traces. We examine how reasoning strategies, prompt construction, and additional modeling considerations affect forecasting performance, and investigate the following research questions:
@@
RQ1: How does TimeReasoner perform compared to state-of-the-art baselines? We evaluate the forecasting accuracy of TimeReasoner across multiple benchmarks to establish its overall effectiveness relative to existing methods.
@@
RQ4: How does TimeReasoner perform under different TSF task configurations? We evaluate the model’s forecasting accuracy and reasoning behavior across varying look-back windows and prediction lengths to understand its adaptability to different temporal settings.
@@
RQ6: Do reasoning traces from TimeReasoner provide explainable insights into its predictions? We examine intermediate reasoning steps and representative case studies to assess their coherence, consistency with observed input patterns, and potential to enhance the interpretability of the model’s forecasts.
Figure 2. Overall framework of TimeReasoner for training-free reasoning-based time series forecasting.:
@@ 2.1. Time Series Forecasting
Time series forecasting has been extensively studied over the past decades (cheng2025comprehensive). Early statistical methods, such as ARIMA (box1968some; zhang2003time), relied on auto-regressive and moving-average formulations with strong theoretical guarantees, but often assumed ideal data properties that rarely hold in practice (hewamalage2021recurrent). With the growth of data and computing power, deep learning-based sequence-to-sequence models have emerged (salinas2020deepar). Recurrent neural networks were among the first to capture temporal dependencies (wen2017multi), but suffered from limited receptive fields and error accumulation in recursive inference (lai2018modeling). Subsequent advances introduced architectures for long-range dependency modeling (zhou2021informer), including self-attention and convolutional designs (li2019enhancing; liu2022scinet). Recent work has also integrated classical techniques—such as trend-seasonal decomposition and time-frequency transforms (zhou2022fedformer; liu2024generative)—into neural networks (wen2020fast; zhou2022film; cheng2025convtimenet). Notably, even simple linear models enhanced with decomposition can be competitive (zeng2023transformers), and slice-based approaches achieve strong performance in long-term forecasting (nie2022time; zhang2023crossformer; cheng2025instructime).
In recent years, large language models (LLMs) have attracted considerable attention for their ability to understand and generate human-like text (kumar2024large; yao2024survey). Beyond their success in natural language processing (zhang2024large), LLMs have shown strong potential in time series analysis (jin2024position). Their application to time series generally follows two paradigms: fine-tuning and prompt-based zero-shot learning (jiang2024empowering). In fine-tuning, a pre-trained LLM is further adapted to time series tasks by training on task-specific data (chang2023llm4ts; chang2024align), enabling it to capture domain patterns and improve performance (liu2025calf; jin2023time), albeit with substantial labeled data and computational costs (cheng2025comprehensive). Prompt-based zero-shot methods instead guide predictions via well-designed prompts without task-specific training (gruver2023large; liu2024lstprompt), offering greater flexibility and efficiency (tang2025time) but sometimes underperforming fine-tuned models in specialized scenarios (wang2024tabletime). Both paradigms underscore the rising interest in LLMs for time series, while revealing challenges in optimizing their effectiveness for such tasks.
@@ 2.2. LLM-based Reasoning
3. Preliminaries
3.1. Problem Definition
Given a multivariate time series $\mathcal{Y}=\{\mathbf{y}_{t}\}_{t=1}^{T}$, where $\mathbf{y}_{t}\in\mathbb{R}^{d}$ denotes the $d$-dimensional target variables at time step $t$, the goal of multivariate time series forecasting (TSF) is to predict future values $\{\mathbf{y}_{t+1},\ldots,\mathbf{y}_{t+H}\}$ over a fixed horizon $H$. At each forecasting step $t$, the model is provided with a lookback window of length $L$, i.e., $\{\mathbf{y}_{t-L},\ldots,\mathbf{y}_{t-1}\}$, summarizing recent historical observations. In addition, the model may access a context vector $\mathcal{Z}\in\mathbb{R}^{k}$ containing supplementary information such as temporal covariates, static attributes, or other relevant features to aid forecasting.
3.2. Motivation of This Work
TSF methods range from statistical models (cheng2025comprehensive), which rely on assumptions such as linearity and stationarity, to data-driven models (lim2021time; benidis2022deep), which learn patterns directly from data. While classical methods like ARIMA (zhang2003time) and Holt-Winters (chatfield1978holt) are interpretable but struggle with complex, context-rich series, deep learning models such as RNNs, TCNs, and Transformers (nie2022time; liu2023itransformer) capture dependencies well but often act as black boxes without explicit reasoning. Recently, LLMs have been explored for TSF, leveraging their ability to process multimodal inputs and structured descriptions (jin2023time), opening opportunities for reasoning-driven forecasting.
Most existing methods still follow a fast-thinking paradigm, generating forecasts in a single forward pass without explicit reasoning over temporal context. However, real-world forecasting often requires step-by-step reasoning over historical trends and auxiliary signals to handle latent dynamics and external factors. These needs, together with the demand for interpretability in high-stakes domains, motivate reframing TSF as a conditional reasoning task guided by structured prompts. The emergence of slow-thinking LLMs, with strong multi-step reasoning capabilities, makes this paradigm increasingly feasible. Building on this insight, we propose TimeReasoner to systematically explore this direction.
@@ 4.1. Overview of TimeReasoner
Figure 2 illustrates the overall architecture of TimeReasoner, a training-free, inference-time reasoning framework for time series forecasting with slow-thinking LLMs. The framework starts with Hybrid Instruction, where instructive prompts are integrated with contextual features (e.g., timestamps, sequential values) to construct structured inputs. These inputs are processed in the Inference-Time Reasoning stage, where slow-thinking LLMs leverage semantic understanding, domain knowledge, and multi-step reasoning capabilities to generate forecasts without task-specific fine-tuning. To further improve forecasting accuracy, Reasoning Strategies—including one-shot multi-step, decoupled multi-step, and rollout multi-step reasoning—are employed, along with uncertainty exploration and inference-time parameter control. Finally, Reasoning Exploration records the reasoning trajectories and supports iterative refinement, backtracking, reflection, and explanation, thereby producing accurate predictions accompanied by interpretable reasoning traces.
4.2. Hybrid Instructions
@@
  • Contextual feature. TimeReasoner is provided with natural language descriptions that offer contextual information like domain-level knowledge and channel-specific semantics. These help align numerical data with real-world meaning, enhancing interpretability and task relevance (williams2024context).
@@ 4.3. Reasoning Strategies
  • One-Shot Reasoning. TimeReasoner performs a single but comprehensive reasoning pass on the given input. During the generation, it processes all necessary information and internally executes the multi-step logic required. The model then outputs the complete set of predicted results, without any subsequent iteration.
  • Decoupled Reasoning. Unlike one-shot reasoning, its process is decoupled: the model generates a partial thought, reflects on it, and then continues, forming a generate–reflect–generate loop within a single inference, enabling refined multi-step reasoning.
@@
4.4. Reasoning Exploration
Beyond generating forecasts, TimeReasoner emphasizes the exploration and analysis of reasoning traces produced by slow-thinking LLMs. This module captures the complete sequence of intermediate reasoning steps during inference, enabling systematic evaluation of their coherence, alignment with input signals, and consistency across different reasoning strategies. The recorded reasoning trajectories can be examined to identify potential errors, detect shifts in contextual patterns, and assess the contribution of specific reasoning steps to the final prediction. Moreover, by supporting iterative refinement, backtracking, and uncertainty estimation, reasoning exploration offers a new perspective for interpreting model behavior and improving forecast reliability. This capability not only enhances transparency in high-stakes applications but also provides a valuable foundation for future integration of human-in-the-loop and self-reflective learning paradigms.
5. Experimental Setup
@@ 5.1. Datasets
To evaluate the generalization of our approach across diverse scenarios, we deliberately select datasets from multiple domains rather than limiting to a few popular benchmarks. The datasets include: ETT (electricity load, 2016–2018)(zhou2021informer); Exchange (foreign exchange rates, 1990–2016)(lai2018modeling); Wind (wind measurements, 2020–2021)(lai2018modeling); AQ (air quality, four years)(zhang2017cautionary); NASDAQ (stock market prices, volumes, and high–low values)(feng2019temporal); and VitalDB (physiological signals sampled every 3 seconds)(cheng2024hmf). These datasets span energy, finance, climate, environmental monitoring, and healthcare, offering heterogeneous temporal patterns, sampling rates, and contextual features. Such diversity supports robust reasoning-based forecasting evaluation beyond narrow benchmark settings and reduces bias toward any single data distribution. Following standard practice, each dataset is chronologically split into training, validation, and testing sets. Table 1 provides detailed descriptions.

### v4

@@ Abstract.
Time series forecasting (TSF) traditionally relies on fast-thinking paradigms that map historical observations directly to future sequences of continuous values. While effective, such approaches often frame forecasting as a pattern-matching problem and tend to overlook explicit reasoning over temporal dynamics and contextual factors, which are critical for modeling long-range dependencies and non-stationary behaviors in real-world scenarios. Recent slow-thinking large language models (LLMs), such as OpenAI o1 and DeepSeek-R1, demonstrate strong inference-time multi-step reasoning abilities. This raises a fundamental question: can slow-thinking LLMs reason over temporal dynamics to support accurate TSF, even without task-specific training? To investigate this question, we present TimeReasoner, a systematic empirical study that reformulates TSF as a conditional reasoning process performed entirely at inference time. TimeReasoner integrates hybrid instructions consisting of task directives, timestamps, sequential values, and optional contextual features, and induces multi-step temporal reasoning in pretrained slow-thinking LLMs through chain-of-thought prompting and rollout-based reasoning strategies. Extensive experiments across diverse TSF benchmarks show that slow-thinking LLMs consistently outperform prior baselines or achieve competitive training-free forecasting performance. Beyond accuracy, we analyze how different inference-time reasoning strategies influence forecasting behaviors, highlighting both the potential and limitations of slow-thinking paradigms for TSF. [^1] https://github.com/realwangjiahao/TimeReasoner
@@ 1. Introduction
Time series forecasting (TSF) is a fundamental data mining task with wide-ranging applications in domains such as finance (tay2001application), energy (deb2017review), and healthcare (kaushik2020ai) where accurate forecasts are critical for strategic planning and effective decision-making under uncertainty (box1968some; pan2025onecast). The goal of TSF is to predict future values of target variables by leveraging historical observations together with contextual features, including temporal covariates and static attributes (hu2016building; wang2024tabletime; jintaozhang).
@@
Over the past decades, numerous approaches have been proposed for time series forecasting (TSF) (cheng2025comprehensive; NEURIPS2023_2e19dab9), which can be broadly categorized into traditional statistical modeling approaches (lim2021time) and data-driven approaches. Traditional statistical models, such as ARIMA (zhang2003time) and Holt–Winters (chatfield1978holt), rely on strong assumptions of linearity, stationarity, and seasonality. While offering clear explainability through explicit model structures, they often struggle to capture complex, nonlinear, or non-stationary temporal patterns. In contrast, data-driven approaches, particularly deep learning–based models, capture intricate temporal patterns such as trends and seasonality through nonlinear architectures, achieving strong performance across standard benchmarks (benidis2022deep; zhang2024trajectory) and demonstrating further improvements in various TSF scenarios (NEURIPS2024_053ee34c; ilbert2024analysing). With recent advances in large language models (LLMs), TSF has also been explored from multimodal understanding and generative modeling perspectives, leveraging the ability of LLMs to incorporate heterogeneous contextual features (jiang2024empowering). Subsequent studies extend this line of work through large-scale LLM integration and generation-oriented architectures (zhang2024large). Despite these advances, most existing LLM-based TSF approaches still treat forecasting as a direct sequence mapping problem, producing predictions in a single inference pass without explicit intermediate reasoning. Consequently, the potential of LLMs for step-wise temporal reasoning—an ability that may be critical for long-horizon and context-aware forecasting—remains largely under-explored (tan2024language; gat2023faithful).
@@
Building on this observation, we further abstract existing TSF methods under a unified perspective and note that most of them largely follow a fast-thinking paradigm (jha2016comparative). Specifically, forecasting is typically performed through primarily single-step inference, mapping historical inputs directly to future targets with limited explicit intermediate reasoning (chow2024towards). Under this paradigm, models often capture short-term temporal dependencies effectively, while systematically integrating heterogeneous contextual information in a step-by-step manner remains challenging (kauppinen2007modeling). In contrast, slow-thinking LLMs, such as OpenAI o1 (o1) and DeepSeek-R1 (guo2025deepseek), demonstrate strong multi-step reasoning abilities in domains such as mathematics, typically by generating explicit intermediate reasoning steps before producing final answers (muennighoff2025s1). Such reasoning processes may offer advantages for modeling long-range dependencies and contextual regime shifts in TSF. However, it remains under-explored whether these reasoning capabilities can be effectively leveraged for time series forecasting, particularly in training-free settings without task-specific fine-tuning.
@@
These questions, centered on the potential of slow-thinking LLMs for TSF, motivate a re-examination of how forecasting problems are formulated and solved. In this work, we investigate whether slow-thinking LLMs can function as forecasters through explicit, inference-time step-by-step reasoning. Specifically, we reformulate TSF as a training-free conditional reasoning task, in which predictions are guided by historical observations together with various contextual features. Accordingly, we propose TimeReasoner, a inference-time reasoning framework that leverages LLMs’ semantic reasoning capabilities to produce both interpretable forecasts and structured reasoning traces. We further analyze how different reasoning strategies, prompt construction, and modeling choices influence forecasting performance, and investigate the following research questions:
@@
RQ1: How does TimeReasoner perform compared to state-of-the-art baselines? We evaluate the forecasting performance of TimeReasoner across multiple benchmarks to establish its overall effectiveness relative to existing methods.
@@
RQ4: How does TimeReasoner perform under different TSF task settings? We evaluate the model’s forecasting accuracy and reasoning trace across varying look-back windows and prediction lengths to understand its adaptability to different temporal settings.
@@
RQ6: Do reasoning traces from TimeReasoner provide explainable insights into its predictions? We examine intermediate reasoning steps and representative case studies to assess their consistency with observed input patterns, and potential to enhance the explaination of the model’s forecasts.
Figure 1. Overall framework of TimeReasoner for training-free reasoning-based time series forecasting.
Overall framework of TimeReasoner for training-free reasoning-based time series forecasting.
@@ 2.1. Time Series Forecasting
Time series forecasting has been extensively studied over the past decades (cheng2025comprehensive). Early statistical methods, such as ARIMA (box1968some; zhang2003time), relied on auto-regressive and moving-average formulations with strong theoretical guarantees, but often assumed ideal data properties that rarely hold in practice (hewamalage2021recurrent). With the growth of data and computing power, deep learning-based sequence-to-sequence models have emerged (salinas2020deepar). Recurrent neural networks were among the first to capture temporal dependencies (wen2017multi), but suffered from limited receptive fields and error accumulation in recursive inference (lai2018modeling). Subsequent advances introduced architectures for long-range dependency modeling (zhou2021informer), including self-attention and convolutional designs (li2019enhancing; liu2022scinet). Recent work has also integrated classical techniques—such as trend-seasonal decomposition and time-frequency transforms (zhou2022fedformer; liu2024generative)—into neural networks (wen2020fast; zhou2022film; zhang2025alphacast). Notably, even simple linear models enhanced with decomposition can be competitive (zeng2023transformers), and slice-based approaches achieve strong performance in long-term forecasting (nie2022time; zhang2023crossformer; cheng2025instructime).
In recent years, large language models (LLMs) have attracted considerable attention for their ability to understand and generate human-like text (kumar2024large; yao2024survey; tao2024hierarchical). Beyond their success in natural language processing (zhang2024large; tao2025values), LLMs have shown strong potential in time series analysis (jin2024position). Their application to time series generally follows two paradigms: fine-tuning and prompt-based zero-shot learning (jiang2024empowering). In fine-tuning, a pre-trained LLM is further adapted to time series tasks by training on task-specific data (chang2023llm4ts; chang2024align), enabling it to capture domain patterns and improve performance (liu2025calf; jin2023time), albeit with substantial labeled data and computational costs (cheng2025comprehensive). Prompt-based zero-shot methods instead guide predictions via well-designed prompts without task-specific training (gruver2023large; liu2024lstprompt), offering greater flexibility and efficiency (tang2025time) but sometimes underperforming fine-tuned models in specialized scenarios (wang2024tabletime). Both paradigms underscore the rising interest in LLMs for time series, while revealing challenges in optimizing their effectiveness for such tasks.
@@ 2.2. LLM-based Reasoning
3. Problem Definition
Given a multivariate time series $\mathcal{Y}=\{\mathbf{y}_{t}\}_{t=1}^{T}$, where $\mathbf{y}_{t}\in\mathbb{R}^{d}$ denotes the $d$-dimensional target variables at time step $t$, the goal of multivariate time series forecasting (TSF) is to predict future values $\{\mathbf{y}_{t+1},\ldots,\mathbf{y}_{t+H}\}$ over a fixed predict window $H$. At each forecasting step $t$, the model is provided with a lookback window of length $L$, i.e., $\{\mathbf{y}_{t-L},\ldots,\mathbf{y}_{t-1}\}$, summarizing recent historical observations. In addition, the model may access a context vector $\mathcal{Z}\in\mathbb{R}^{k}$ containing supplementary information such as temporal covariates, static attributes, or other relevant features to aid forecasting.
@@ 4. The Proposed TimeReasoner
In this section, we present the overall architecture of TimeReasoner, detailing how it leverages LLMs to perform inference-time reasoning for time series forecasting.
@@ 4.1. Overview of TimeReasoner
Figure 1 illustrates the overall architecture of TimeReasoner, a training-free, inference-time reasoning framework for time series forecasting using slow-thinking LLMs. The framework starts with Hybrid Instruction, where instructive prompts are integrated with contextual features (e.g., timestamps, sequential values) to construct structured inputs. These inputs are processed in the Inference-Time Reasoning stage, where slow-thinking LLMs leverage semantic understanding, domain knowledge, and multi-step reasoning capabilities to generate forecasts without task-specific fine-tuning. To further improve forecasting accuracy, reasoning strategies including one-shot multi-step, decoupled multi-step, and rollout multi-step reasoning are employed, along with uncertainty exploration and inference-time parameter control. Finally, Reasoning Exploration records the reasoning trajectories and supports iterative refinement, backtracking, reflection, and explanation, thereby producing accurate predictions accompanied by interpretable reasoning traces.
4.2. Hybrid Instruction
@@
  • Contextual features. TimeReasoner is provided with natural language descriptions that offer contextual information like domain-level knowledge and channel-specific semantics. These help align numerical data with real-world meaning, enhancing explainability and task relevance (williams2024context).
@@
4.3. Inference-Time Reasoning
Based on the constructed multimodal prompts, TimeReasoner performs inference-time reasoning by leveraging the semantic understanding and complex reasoning capabilities of slow-thinking LLMs. Specifically, the model first interprets the forecasting objective and contextual descriptions to establish a high-level semantic grounding of the task, including the prediction horizon, domain constraints, and plausible temporal behaviors. It then reasons over the raw time series values and timestamps jointly, identifying salient temporal patterns such as trends, periodicity, regime shifts, and potential anomalies directly in the original value space. During generation, the LLM incrementally integrates these observations, reflecting on recent dynamics and historical regularities to form a coherent forecasting trajectory. This inference-time reasoning process allows TimeReasoner to adaptively combine numerical evidence with domain semantics, enabling flexible and interpretable time series forecasting without task-specific parameter tuning.
@@ 4.4. Reasoning Strategies
  • One-Shot Reasoning. TimeReasoner performs a single, comprehensive reasoning pass on the given input. During the generation, it processes all necessary information and internally executes the multi-step logic required. The model then outputs the complete set of predicted results, without any subsequent iteration.
  • Decoupled Reasoning. Unlike one-shot reasoning, the reasoning process is decoupled: the model generates a partial thought, reflects on it, and then continues, forming a generate–reflect–generate loop within a single inference, enabling multi-step reasoning.
@@
4.5. Reasoning Exploration and Trajectory
Beyond directly generating numerical forecasts, TimeReasoner leverages the slow-thinking capability of LLMs to perform iterative reasoning over time series dynamics at inference time. Given historical observations and timestamps, the model incrementally analyzes recent temporal behaviors, such as local trends, periodic patterns, and potential regime changes, and reflects on how these dynamics may evolve over the predict window. During this process, the LLM may revisit earlier hypotheses, backtrack from inconsistent assumptions, and refine its interpretation of the time series when new evidence (e.g., abrupt fluctuations or weakened seasonality) is identified. This iterative reasoning and reflection enable the model to form a coherent forecasting rationale rather than relying on a single-pass mapping from past to future values.
To obtain a stable and reliable prediction, we execute the slow-thinking forecasting procedure multiple times under the same input and aggregate the generated results; specifically, we perform three independent generations and take the mean of the resulting forecast sequences as the final output. This aggregation strategy mitigates variability introduced by stochastic inference and provides a more robust estimate of future time series values, while preserving the explainability afforded by explicit reasoning traces.
5. Experimental Settings
In this section, we describe the experimental settings for evaluating TimeReasoner, covering datasets, baselines, implementation details.
@@ 5.1. Datasets
To evaluate the generalization of our approach across diverse scenarios, we deliberately select datasets from multiple domains rather than a few popular benchmarks. The datasets include: ETT (electricity load, 2016–2018) (zhou2021informer); Exchange (foreign exchange rates, 1990–2016) (lai2018modeling); Wind (wind measurements, 2020–2021) (lai2018modeling); AQ (air quality, four years) (zhang2017cautionary); NASDAQ (stock market prices, volumes, and high–low values) (feng2019temporal); and VitalDB (physiological signals sampled every 3 seconds) (cheng2024hmf). These datasets span energy, finance, climate, environmental monitoring, and healthcare, offering heterogeneous temporal patterns, sampling rates, and contextual features. Such diversity supports robust reasoning-based forecasting evaluation beyond narrow benchmark settings and reduces bias toward any single data distribution. Following standard practice, each dataset is chronologically split into training, validation, and testing sets. Table 1 provides detailed descriptions.